Hi All, Please help me to understand the below problem, differentiate cos ​-1 (1-x/1+x) w.r.t x ?

Hi All,
Please help me to understand the below problem,
differentiate cos​-1(1-x/1+x) w.r.t x ?
 

Grade:12th pass

1 Answers

Vinod Prajapati
6 years ago
Substitute x=tan^2theta therefore theta=tan^-1(√x). So,d/dx[cos^​-1(1-tan^theta/1+tan^2theta)]. =d/dX[Cos^-1(cos2theta)] =d/dX[2*tan^-1(√x)]. Ans=2[1/1+x]*1/2√xSo the answer is 1/√x(1+x)